OpenStudy (bbypanda16):

Please helppp Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

type this equation into a site I found awhile back, go to mathway.com. It can solve almost any equation

OpenStudy (bbypanda16):

@j-bird Okay, thank you :)

OpenStudy (texaschic101):

| x - 2 | + 10 = 12 | x - 2 | = 12 - 10 | x - 2 | = 2 x - 2 = 2 x = 2 + 2 x = 4 -(x - 2) = 2 -x + 2 = 2 -x = 2 - 2 -x = 0 x = 0 answers are : x = 0 and x = 4

OpenStudy (bbypanda16):

@texaschic101 Thank you :)

OpenStudy (texaschic101):

sure thing :)